# Joint Selective State Space Model and Detrending for Robust Time Series Anomaly Detection

### v1

@@ I Introduction
To capture long-term dependencies, various sequence models, including Temporal Convolusional Networks (TCNs) [15], Recurrent Neural Networks (RNNs) [8, 9] and Transformers [11, 12], have been explored for TSAD. However, existing methods still encounter difficulties due to their intrinsic characteristics, such as limited context window, high memory costs, or unstable gradient flow. Recent research on the selective state space model (S6) [16] has addressed the shortcomings of the aforementioned sequence models and demonstrated its excellent long-term modeling capabilities in other domains. Moreover, the selective nature of the S6 model allows it to discard abnormal information and generate reliable reconstructed output for TSAD. Despite offering a potential solution to model long-range dependencies, there has been no research exploring its application in TSAD.
The generalization issue for non-stationary data stems from trends in non-stationary data. The erractic distribution of trends can result in significant fluctuations in data magnitude, leading to erroneously high reconstruction errors in regions with previously unseen trend patterns in the training set, ultimately causing false alarms [14]. Traditional time series decomposition methods such as Seasonal-Trend decomposition using LOESS (STL) [17] and Hodrick–Prescott (HP) trend filter [18] are frequently used to mitigate the impact of these trends. However, achieving optimal decomposed results may not yield the best detection performance. To address this, several approaches have tried to integrate decomposition methods into detectors for joint optimization. Nevertheless, most of them rely on Moving Average (MA) with a fixed kernel size [19, 20, 14], lacking the flexibility for broader scenarios.
@@ III-B S6 Model
$\bm{h}_{t,d}$ $=\overline{\bm{A}}_{t,d}\odot\bm{h}_{t-1,d}+\overline{\bm{B}}_{t,d}\cdot{x}_{t,d}$ (1)
@@ IV-A HP Trend Filter
The input data was initially segmented into subsequences using sliding windows of length $W$. Each input window $\bm{x}^{\rm in}\in\mathbb{R}^{W\times D}$ underwent detrending using the HP trend filter to remove prominent trend components and ensure stable input for subsequent processing. The HP-based trend component $\bm{\tau}_{\rm HP}$ was computed via equation (3), while the HP-based seasonality was derived as $\bm{s}_{\rm HP}=\bm{x}^{\rm in}-\bm{\tau}_{\rm HP}$.
@@ IV-B DMamba Blocks
The HP-based seasonality $\bm{s}_{\rm HP}$ was first projected to a $D_{m}$-dimensional space by an embedding network ${\rm Emb(\cdot)}$, and then it underwent $L$ DMamba blocks. Each block comprised a Mamba block [16] and an Adaptive MA (AMA) module. The input of each block was the output of the previous block, denoted as $\bm{x}^{l}=\bm{y}^{l-1}$ with $\bm{y}^{0}={\rm Emb}(\bm{s}_{\rm HP})$. For clarity, the superscript $l$ was omitted in the following content. The input $\bm{x}$ first entered the Mamba block shown in Fig. 3 and could be formulated by equation (4):
@@ V-A3 Implementation details
All baselines were implemented based on PyOD [37] or their publicly available codes. We utilized Peak-Over-Threshold (POT) [38] for threshold selection. The smoothness parameter $\lambda$ was set to $10^{4}$. The block size $L$ was $3$, the model size $D_{m}$ was $32$, the state size was $16$ and the window size $W$ was $100$. The training was conducted for $7$ epochs with a batch size of $128$, using the AdamW [39] optimizer with a learning rate of $10^{-4}$. Given that the common point-adjustment strategy [9] could lead to misguided rankings [40], the recently proposed affiliation-based precision (P-AF), recall (R-AF), and F1-score (F1-AF) [42] were adopted as evaluation metrics, as they could offer an intuitive interpretation for time-series data. Since F1-AF represents the harmonic mean of P-AF and R-AF, it was employed to evaluate the overall performance, a common practice in the TSAD community.
@@ V-B Results
To further verify the necessity of the HP trend filter and the AMA mechanism, an ablation analysis was conducted and the results were shown in Tab. III. The results demonstrated that applying only one might not bring improvement. For instance, solely applying the HP trend filter in the NASA dataset or only utilizing AMA in the SWaT dataset led to a performance drop. However, when both were applied, the HP trend filter provided stable input for AMA’s period estimation, where AMA and the detector could symbiotically optimize performance and capture more complex patterns, resulting in the best performance over all datasets.

### v2

@@ I Introduction
Normal behavior in time-series data typically involves long-term dependencies[1]. To capture these dependencies and enhance the modeling of normal behavior, various sequence models including Temporal Convolutional Networks (TCNs) [15], Recurrent Neural Networks (RNNs) [8, 9] and Transformers [11, 12] have been explored for TSAD. However, existing methods still encounter difficulties due to their intrinsic characteristics, such as limited context window, high memory costs, or unstable gradient flow. Recent research on the selective state space model (S6) [16] has addressed the shortcomings of the aforementioned sequence models and demonstrated its excellent long-term modeling capabilities in other domains. Moreover, the selective nature of the S6 model allows it to discard abnormal information and generate reliable reconstructed output for TSAD. Despite offering a potential solution to model long-range dependencies, there has been no research exploring its application in TSAD.
The generalization issue for non-stationary data stems from trends in non-stationary data. The erratic distribution of trends can result in significant fluctuations in data magnitude, leading to erroneously high reconstruction errors in regions with previously unseen trend patterns in the training set, ultimately causing false alarms [14]. Traditional time series decomposition methods such as Seasonal-Trend decomposition using LOESS (STL) [17] and Hodrick–Prescott (HP) trend filter [18] are frequently used to mitigate the impact of these trends. However, achieving optimal decomposed results may not yield the best detection performance. To address this, several approaches have tried to integrate decomposition methods into detectors for joint optimization. Nevertheless, most of them rely on Moving Average (MA) with a fixed kernel size [19, 20, 14], lacking the flexibility for broader scenarios. Other methods rely on DNN models for detrending [21, 22]. Though they allow end-to-end optimization, their dependence on training data reintroduces generalization issues.
@@ III-B S6 Model
$\bm{h}_{t,d}$ $=\overline{\bm{A}}_{t,d}\odot\bm{h}_{t-1,d}+\overline{\bm{B}}_{t,d}\cdot{x}_{t,d},$ (1)
@@ IV-A HP Trend Filter
The input data was initially segmented into subsequences using sliding windows of length $W$. Each input window $\bm{x}^{\rm in}\in\mathbb{R}^{W\times D}$ underwent detrending using HP trend filter. To ensure the extraction of global trends, the current input window was first combined with historical windows, and then the HP-based trend component $\bm{\tau}_{\rm HP}$ was computed via Eq. (3). The HP-based seasonality was further derived as $\bm{s}_{\rm HP}=\bm{x}^{\rm in}-\bm{\tau}_{\rm HP}$.
@@ IV-B DMamba Blocks
The HP-based seasonality $\bm{s}_{\rm HP}$ was first projected to a $D_{m}$-dimensional space by an embedding network ${\rm Emb(\cdot)}$, and then it underwent $L$ DMamba blocks. Each block comprised a Mamba block [16] and an Adaptive MA (AMA) module. The input of each block was the output of the previous block, denoted as $\bm{x}^{l}=\bm{y}^{l-1}$ with $\bm{y}^{0}={\rm Emb}(\bm{s}_{\rm HP})$. For clarity, the superscript $l$ was omitted in the following content. The input $\bm{x}$ first entered the Mamba block shown in Fig. 3 and could be formulated by Eq. (4):
@@ V-A3 Implementation details
All baselines were implemented based on PyOD [39] or their publicly available codes. We utilized Peak-Over-Threshold (POT) [40] for threshold selection. The smoothness parameter $\lambda$ was set to $10^{4}$. The block size $L$ was $3$, the model size $D_{m}$ was $32$, the state size was $16$ and the window size $W$ was $100$. The training was conducted for $7$ epochs with a batch size of $128$, using the AdamW [41] optimizer with a learning rate of $10^{-4}$. Given that point-wise metrics can be influenced by lengthy anomalies and point-adjustment strategy [9] could lead to misguided rankings [42], the recently proposed affiliation-based precision (P-AF), recall (R-AF), and F1-score (F1-AF) [43] were adopted as evaluation metrics. These metrics offer evaluations at the level of abnormal events rather than individual points, making them suitable for TSAD. Since F1-AF represents the harmonic mean of P-AF and R-AF, it was employed to evaluate the overall performance, a common practice in the TSAD community.
@@ V-B Results
To further verify the necessity of the HP trend filter and the AMA mechanism, an ablation analysis was conducted and the results were shown in Tab. III. The results demonstrated that applying only one might not bring improvement. For instance, solely applying the HP trend filter in the NASA dataset or only utilizing AMA in the SWaT dataset led to a performance drop. However, when both were applied, the HP trend filter provided stable input for AMA’s period estimation, where AMA and the detector could symbiotically optimize performance and capture more complex patterns, resulting in the best performance over all datasets. Additionally, a visualization in Fig. 4 clearly showcased the effectiveness of the proposed detrending mechanisms.
Figure 4: A visualization from the SWaT dataset, which contains non-stationary data. The first row shows the input signal and extracted trend. The second row displays anomaly scores from various detectors. Baseline methods were influenced by trends, leading to false alarms, while our proposed method generated reliable scores by incorporating detrending mechanisms.